# A Finer View of the Parameterized Landscape of Labeled Graph Contractions

### v1

@@ 3.1.3 Updating Dynamic Programming Tables
$(\Leftarrow)$ Assume the three conditions of the lemma are met. By condition (ii), there exists a valid signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ at the child node $t^{\prime}$, which means there is a valid partial solution $P^{\prime}=(\sigma^{\prime}_{p},\tau^{\prime}_{p},\mathcal{R}^{\prime}_{p},\mathcal{A}^{\prime}_{p})$ for the subgraph $G_{t^{\prime}}$ that corresponds to $S^{\prime}$. We construct a new partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for the subgraph $G_{t}$ by extending $P^{\prime}$ to include vertex $x$. For all $u\in V(G_{t^{\prime}})$, we set $\sigma_{p}(u)=\sigma^{\prime}_{p}(u)$ and $\tau_{p}(u)=\tau^{\prime}_{p}(u)$. For the new vertex $x\notin V(H)$, its parent must lie outside the current subtree. Thus, we set $\sigma_{p}(x)=\star$ and $\tau_{p}(x)=\star$, consistent with condition (i). Since no new edges are introduced involving $x$, we set $\mathcal{R}_{p}=\mathcal{R}^{\prime}_{p}$ and $\mathcal{A}_{p}=\mathcal{A}^{\prime}_{p}$, consistent with condition (iii). This constructed tuple $P$ is a valid partial solution for $G_{t}$, and its signature is precisely $S$. Therefore, $\mathbf{d}[S,t]$ is set to True.
$(\Rightarrow)$ Now, assume that $\mathbf{d}[S,t]=\text{true}$. This means there exists a valid partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for $G_{t}$ whose signature is $S$. We can define a partial solution $P^{\prime}$ for the smaller subgraph $G_{t^{\prime}}$ by simply restricting $P$. For all $u\in V(G_{t^{\prime}})$, we set $\sigma^{\prime}_{p}(u)=\sigma_{p}(u)$, $\tau^{\prime}_{p}(u)=\tau_{p}(u)$, $\mathcal{R}^{\prime}_{p}=\mathcal{R}_{p}$, and $\mathcal{A}^{\prime}_{p}=\mathcal{A}_{p}$. This new tuple $P^{\prime}$ is a valid partial solution for $G_{t^{\prime}}$, and its signature $S^{\prime}$ must exist in the DP table, meaning $\mathbf{d}[S^{\prime},t^{\prime}]=\text{true}$. This proves condition (ii). Since $x\notin V(H)$ and is a new, isolated vertex, it cannot be a parent to any other vertex in the subtree. This requires that in any valid partial solution, $\sigma_{p}(x)=\star$ and $\tau_{p}(x)=\star$, which proves condition (i). Finally, because no new edges or relationships were formed, the sets of requester pairs and pseudo-adjacencies remain unchanged from $t^{\prime}$ to $t$. Thus, $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$, which proves condition (iii). ∎
@@
Let $t$ be an introduce vertex node with child $t^{\prime}$, and let $X_{t}=X_{t^{\prime}}\cup\{x\}$ and $x\in V(H)$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$. Assign the value of $\mathbf{d}[S,t]=\texttt{True}$ if and only if the following conditions are met:
  1. (i)
    $\sigma(x)=x$ and $\tau(x)=x$,
  2. (ii)
    there exists a signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})\in I(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$,
  3. (iii)
    $\mathcal{A}=\mathcal{A}^{\prime}$, and every element in $\mathcal{R}\setminus\mathcal{R}^{\prime}$ is of the form
    • $(x,u,\textsf{unsat})$ for some $u\in X_{t}$ such that $x$ and $\tau(u)$ are adjacent in $H$, or
    • $(x,u,\textsf{sat})$ for some $u\in X_{t}$ such that $x$ and $\tau(u)$ are not adjacent in $H$.
    The above condition implies that the only tuple introduced in $\mathcal{R}$ contains $x$ in it.
@@
$(\Leftarrow)$ Assume the three conditions of the lemma are met. By condition (ii), there exists a valid signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ at the child node $t^{\prime}$, which means there is a valid partial solution $P^{\prime}=(\sigma^{\prime}_{p},\tau^{\prime}_{p},\mathcal{R}^{\prime}_{p},\mathcal{A}^{\prime}_{p})$ for the subgraph $G_{t^{\prime}}$. We construct a new partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for the subgraph $G_{t}$ by extending $P^{\prime}$ to include vertex $x$. For all $u\in V(G_{t^{\prime}})$, we set $\sigma_{p}(u)=\sigma^{\prime}_{p}(u)$ and $\tau_{p}(u)=\tau^{\prime}_{p}(u)$. For the new vertex $x\in V(H)$, it must be the root of its own witness set. Thus, we set $\sigma_{p}(x)=x$ and $\tau_{p}(x)=x$, which is consistent with condition (i). The pseudo-adjacency set is unchanged as no new edges from $G$ are introduced, so we set $\mathcal{A}_{p}=\mathcal{A}^{\prime}_{p}$, consistent with condition (iii). As $x$ is now a requester vertex, its relationships with other requesters in the bag must be recorded. Based on condition (iii), we define $\mathcal{R}_{p}=\mathcal{R}^{\prime}_{p}\cup\{(x,u,\textsf{unsat})\mid(x,\tau(u))\in E(H)\}\cup\{(x,u,\textsf{sat})\mid(x,\tau(u))\notin E(H)\}$. The constructed tuple $P$ is a valid partial solution for $G_{t}$, and its signature is precisely $S$. Therefore, $\mathbf{d}[S,t]$ is set to True.
$(\Rightarrow)$ Now, assume that $\mathbf{d}[S,t]=\text{true}$. This means there exists a valid partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for $G_{t}$ whose signature is $S$. We can define a partial solution $P^{\prime}$ for the smaller subgraph $G_{t^{\prime}}$ by simply restricting $P$. For all $u\in V(G_{t^{\prime}})$, we set $\sigma^{\prime}_{p}(u)=\sigma_{p}(u)$ and $\tau^{\prime}_{p}(u)=\tau_{p}(u)$. The sets $\mathcal{R}^{\prime}_{p}$ and $\mathcal{A}^{\prime}_{p}$ are simply the projections of $\mathcal{R}_{p}$ and $\mathcal{A}_{p}$ onto pairs of vertices in $X_{t^{\prime}}$. The resulting tuple $P^{\prime}$ is a valid partial solution for $G_{t^{\prime}}$ since it omits the new vertex $x$. Its signature $S^{\prime}$ must exist in the DP table, meaning $\mathbf{d}[S^{\prime},t^{\prime}]=\text{true}$. This proves condition (ii). Since $x\in V(H)$, by the properties of a valid witness structure, $x$ must be the root of its witness tree and its own originator. This means $\sigma(x)=x$ and $\tau(x)=x$, which proves condition (i). Finally, as no edges from $G$ are introduced, the pseudo-adjacency set for pairs in $X_{t}$ remains unchanged from $t^{\prime}$ to $t$. Any new entries in the requester set $\mathcal{R}$ must involve the new vertex $x$ and must be correctly classified based on its relationships to other requesters, as captured in condition (iii). ∎
Forget Vertex Node.
Let $t$ be a forget vertex node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}\setminus\{x\}$ for some vertex $x\in X_{t^{\prime}}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$, and let $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ be a signature at node $t^{\prime}$.
Definition 3.13.
We say that $S$ and $S^{\prime}$ are forget-vertex compatible if the following conditions hold:
  • (C1) The vertex $x$ is not an explorer with respect to $\sigma^{\prime}$, as per Definition 3.7.
  • (C2) For every $y\in X_{t^{\prime}}\setminus\{x\}$, we have $\sigma(y)=\sigma^{\prime}(y)$ and $\tau(y)=\tau^{\prime}(y)$.
  • (C3) $\mathcal{R}$ is the restriction of $\mathcal{R}^{\prime}$ to pairs not involving $x$, and $\mathcal{A}$ is the restriction of $\mathcal{A}^{\prime}$ to pairs not involving $x$.
Let $F(S)$ denote the set of all signatures at $t^{\prime}$ that are forget-vertex compatible with $S$.
@@
Let $t$ be a forget vertex node with child $t^{\prime}$, and let $X_{t}=X_{t^{\prime}}\setminus\{x\}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$. Then, $\mathbf{d}[S,t]=\texttt{True}$ if and only if the following conditions hold:
  1. (i)
    There exists a signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})\in F(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$.
  2. (ii)
    The requester set $\mathcal{R}^{\prime}$ contains no entry of the form $(x,u,\textsf{unsat})$ for any $u\in X_{t^{\prime}}\setminus\{x\}$.
@@
$(\Leftarrow)$ Assume the two conditions of the lemma are met. By condition (i), there exists a valid signature $S^{\prime}$ at $t^{\prime}$ corresponding to a valid partial solution $P^{\prime}$ for the subgraph $G_{t^{\prime}}$. We construct a new partial solution $P$ for $G_{t}$ by restricting $P^{\prime}$ to $V(G)\setminus\{x\}$. Since $P^{\prime}$ is valid, all adjacency requirements imposed on $G_{t^{\prime}}$ must hold. Condition (ii) explicitly guarantees that the vertex $x$ does not have any unsat adjacency requirements with any other vertex $u\in X_{t}$. Because $x$ is not an explorer (C1), all required parent/originator relations for $x$ must be contained within $G_{t^{\prime}}$. By C3, the signature $S$ is precisely the valid signature $S^{\prime}$ with all information about $x$ correctly removed. Since all constraints involving $x$ are resolved within $G_{t^{\prime}}$ and $P^{\prime}$ is valid, the restriction $P$ is a valid partial solution for $G_{t}$. Therefore, $\mathbf{d}[S,t]$ is set to True.
$(\Rightarrow)$ Now, assume that $\mathbf{d}[S,t]=\text{true}$. This means there exists a valid partial solution $P$ for $G_{t}$ whose signature is $S$. This solution $P$ must be the restriction of some valid partial solution $P^{\prime}$ for $G_{t^{\prime}}$. This proves the existence of a compatible signature $S^{\prime}$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\text{true}$, proving condition (i). For $P^{\prime}$ to be a valid partial solution, all requirements within $G_{t^{\prime}}$ must be satisfied. If $\mathcal{R}^{\prime}$ contained an entry $(x,u,\textsf{unsat})$, the required edge in $H$ between $\tau(x)$ and $\tau(u)$ would not have been realized by the edges in $G_{t^{\prime}}$, contradicting the validity of $P^{\prime}$. Thus, no unsat entry involving $x$ can exist in $\mathcal{R}^{\prime}$. This proves condition (ii). ∎
Introduce Edge Node
Let $t$ be an introduce edge node with child $t^{\prime}$, so that $X_{t}=X_{t^{\prime}}$, and let the edge $(u,v)\in E(G)$ be introduced at $t$, with $u,v\in X_{t}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a candidate signature at $t$.
Definition 3.14.
We say that $S$ is introduce-edge compatible with respect to its counterpart at $t^{\prime}$, denoted $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$, if the following hold:
  • (C1) For every $x\in X_{t}\setminus\{u,v\}$, we have $\sigma(x)=\sigma^{\prime}(x)$ and $\tau(x)=\tau^{\prime}(x)$.
  • (C2) At most one of the following statements is true for the values of $\sigma$ and $\tau$ at $u$ and $v$:
    • (C2.1) No change in values. Formally, $\sigma^{\prime}(u)=\sigma(u)$, $\tau^{\prime}(u)=\tau(u)$ and $\sigma^{\prime}(v)=\sigma(v)$, $\tau^{\prime}(v)=\tau(v)$.
    • (C2.2) $v$’s set merges into $u$’s set. Formally, $\sigma(u)=\sigma^{\prime}(u)$, $\tau(u)=\tau^{\prime}(u)$, $\sigma(v)\neq\sigma^{\prime}(v)$, and $\tau(v)=\tau(u)$. This change represents $u$ becoming the parent of $v$ or the root of $v$’s witness tree.
    • (C2.3) $u$’s set merges into $v$’s set. Formally, $\sigma(v)=\sigma^{\prime}(v)$, $\tau(v)=\tau^{\prime}(v)$, $\sigma(u)\neq\sigma^{\prime}(u)$, and $\tau(u)=\tau(v)$.
  • (C3) $\mathcal{R}^{\prime}\setminus\mathcal{R}\subseteq\{(u,v,\textsf{unsat}),(v,u,\textsf{unsat})\}$ and $\mathcal{A}^{\prime}\subseteq\mathcal{A}$.
Let $I_{e}(S)$ denote the set of all signatures at $t^{\prime}$ that are introduce-edge compatible with $S$.
Lemma 3.6.
Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u,v\in V(H)$. A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and the following conditions hold:
  1. (i)
    Condition C2.1 holds.
  2. (ii)
    The pseudo-adjacency set is unchanged: $\mathcal{A}=\mathcal{A}^{\prime}$.
  3. (iii)
    The requester set is updated only for $(u,v)$: $\mathcal{R}=\mathcal{R}^{\prime}\setminus\{(u,v,\textsf{unsat})\}\cup\{(u,v,\textsf{sat})\}$, if $(u,v,\textsf{unsat})\in\mathcal{R}^{\prime}$. Otherwise $\mathcal{R}=\mathcal{R}^{\prime}$.

### v2

@@ 3.1 Algorithmic Result
$(\Rightarrow)$ If $\texttt{dp}[t;S]=\texttt{True}$ with witness structure $\mathcal{W}$, then $\mathcal{W}$ restricted to $G_{t^{\prime}}$ induces a signature $S^{\prime}$. Since $X_{t}=X_{t^{\prime}}$, only the adjacency set $\mathcal{A}_{t}$ can change. If the introduced edge $e=\{u,v\}$ connects different parts $P_{u}$ and $P_{v}$, it witnesses a new pseudo-adjacency $(P_{u},P_{v})\in\mathcal{A}_{t}$, while all other components remain identical to $S^{\prime}$.
$(\Leftarrow)$ Given a witness structure for $G_{t^{\prime}}$ inducing $S^{\prime}$, we add the edge $e=\{u,v\}$ to form $G_{t}$. If $P_{u}\neq P_{v}$, this edge creates a new adjacency between witness sets, which is correctly reflected in $\mathcal{A}_{t}$ by the compatibility condition. Because $S$ is valid, this new adjacency does not violate any non-adjacency constraints in $\mathcal{R}_{t}$, so the structure remains a valid partial $H$-witness structure for $G_{t}$. ∎
Forget Vertex Node
Let $t$ be a forget vertex node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}\setminus\{x\}$ for some vertex $x\in X_{t^{\prime}}$.
Let $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ be a signature at node $t$, and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ be a signature at node $t^{\prime}$. Let $P_{x}\in\mathcal{P}^{\prime}$ be the part containing $x$, and $P_{x}^{*}=P_{x}\setminus\{x\}$.
Definition 3.9 (Forget-Vertex Compatibility).
Valid signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ at node $t$ and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ at child $t^{\prime}$ are forget-vertex compatible if one of the following two cases holds:
  • $x$ is not the only vertex in its part ($P_{x}\neq\{x\}$):
    • (C1)
      $\mathcal{P}=(\mathcal{P}^{\prime}\setminus\{P_{x}\})\cup\{P_{x}^{*}\}$.
    • (C2)
      For all $P\in\mathcal{P}^{\prime}\setminus\{P_{x}\}$, $\tau(P)=\tau^{\prime}(P)$. For the resulting part $P_{x}^{*}$:
      1. 1.
        If $\tau^{\prime}(P_{x})=x$, then $\tau(P_{x}^{*})=\downarrow$.
      2. 2.
        Otherwise, $\tau(P_{x}^{*})=\tau^{\prime}(P_{x})$.
    • (C3)
      $\mathcal{A}_{t}$ and $\mathcal{R}_{t}$ are the restrictions of $\mathcal{A}_{t^{\prime}}$ and $\mathcal{R}_{t^{\prime}}$ to $\mathcal{P}$.
  • $x$ is the last vertex in its part ($P_{x}=\{x\}$):
    • (C1)
      $\mathcal{P}=\mathcal{P}^{\prime}\setminus\{P_{x}\}$.
    • (C2)
      For all $P\in\mathcal{P}$, $\tau(P)=\tau^{\prime}(P)$.
    • (C3)
      $\mathcal{A}_{t}$ and $\mathcal{R}_{t}$ are the restrictions of $\mathcal{A}_{t^{\prime}}$ and $\mathcal{R}_{t^{\prime}}$ to $\mathcal{P}$.
    • (C4)
      For all $P_{j}\in\mathcal{P}$, if $(P_{x},P_{j})\notin\mathcal{R}_{t^{\prime}}$, then $(P_{x},P_{j})\in\mathcal{A}_{t^{\prime}}$.
@@
Let $t$ be a forget vertex node with child $t^{\prime}$. For a valid signature $S$, $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exists a signature $S^{\prime}$ such that $\texttt{dp}[t^{\prime};S^{\prime}]=\texttt{True}$ and $S,S^{\prime}$ are forget-vertex compatible.
@@
$(\Rightarrow)$ If $\texttt{dp}[t;S]=\texttt{True}$ with witness structure $\mathcal{W}$, then $\mathcal{W}$ induces a signature $S^{\prime}$ for $G_{t^{\prime}}$ containing $x$. Since $G_{t}=G_{t^{\prime}}$, the only difference is the restriction of the partition to $X_{t}\setminus\{x\}$. If $x$ was the last vertex of its part in the bag ($P_{x}=\{x\}$), any witness set $W_{j}$ required to be adjacent to $W_{x}$ in $H$ (where $(P_{x},P_{j})\notin\mathcal{R}_{t^{\prime}}$) must have already established an adjacency in $G_{t^{\prime}}$, satisfying (C4).
$(\Leftarrow)$ A valid witness structure for $G_{t^{\prime}}$ remains valid for $G_{t}$. The compatibility cases ensure $S$ correctly tracks the representative status after $x$ is removed. If $P_{x}$ is removed from the bag (Case 2), condition (C4) guarantees that $W_{x}$ has already satisfied all required adjacencies with other witness sets. Since $S$ is valid, the forgotten witness set $W_{x}$ does not violate any global $H$-contraction requirements, so $\texttt{dp}[t;S]=\texttt{True}$. ∎
Join Node
Let $t$ be a join node with children $t_{L}$ and $t_{R}$, such that $X_{t}=X_{t_{L}}=X_{t_{R}}$.
Definition 3.10 (Join-Node Compatibility).
Valid signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ at node $t$, $S_{L}=(\mathcal{P}_{L},\tau_{L},\mathcal{A}_{L},\mathcal{R}_{L})$ and $S_{R}=(\mathcal{P}_{R},\tau_{R},\mathcal{A}_{R},\mathcal{R}_{R})$ at children $t_{L},t_{R}$ are join-compatible if the following conditions hold:
  • (C1)
    $\mathcal{P}=\mathcal{P}_{L}=\mathcal{P}_{R}$. Furthermore, for every $P\in\mathcal{P}$:
    • If $\tau_{L}(P)\in V(H)$ and $\tau_{R}(P)=\uparrow$, then $\tau(P)=\tau_{L}(P)$.
    • If $\tau_{R}(P)\in V(H)$ and $\tau_{L}(P)=\uparrow$, then $\tau(P)=\tau_{R}(P)$.
    • If $\tau_{L}(P)=\downarrow$ or $\tau_{R}(P)=\downarrow$, then $\tau(P)=\downarrow$.
    • Otherwise, $\tau_{L}(P)=\tau_{R}(P)=\tau(P)$.
  • (C2)
    $\mathcal{A}_{t}=\mathcal{A}_{L}\cup\mathcal{A}_{R}$.
  • (C3)
    $\mathcal{R}_{t}=\mathcal{R}_{L}=\mathcal{R}_{R}$.
@@
Let $t$ be a join node with children $t_{L},t_{R}$. For a valid signature $S$, $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exist signatures $S_{L},S_{R}$ such that $\texttt{dp}[t_{L};S_{L}]=\texttt{True}$, $\texttt{dp}[t_{R};S_{R}]=\texttt{True}$, and $S,S_{L},S_{R}$ are join-compatible.
@@
$(\Rightarrow)$ If $\texttt{dp}[t;S]=\texttt{True}$, the witness structure $\mathcal{W}$ for $G_{t}$ restricted to $G_{t_{L}}$ and $G_{t_{R}}$ must induce $S_{L}$ and $S_{R}$. Since $V(G_{t_{L}})\cap V(G_{t_{R}})=X_{t}$, the bag partitions must be identical. The status $\tau(P)$ correctly merges the representative’s state (present, missing, or forgotten) from both branches. Since $E(G_{t})=E(G_{t_{L}})\cup E(G_{t_{R}})$, any set-to-set adjacency witnessed in $G_{t}$ must appear in $\mathcal{A}_{L}\cup\mathcal{A}_{R}$.
$(\Leftarrow)$ Given compatible $S_{L},S_{R}$, we merge their witness structures by taking the union of corresponding witness sets. Compatibility ensures they agree on the bag partition and representatives. The union $\mathcal{A}_{t}=\mathcal{A}_{L}\cup\mathcal{A}_{R}$ accounts for all edges in $G_{t}$. Since $S$ is valid, this union satisfies all non-adjacency requirements in $\mathcal{R}_{t}$, making the merged structure a valid partial $H$-witness structure for $G_{t}$. ∎
We are finally at a position to present proof of algorithm mentioned in Theorem 1.